Can this problem be done by elimination are subsitution
if so can you show me which one and how to do it?
y+4x=-5
y+5x=3

Found 2 solutions by rfer, ewatrrr:
Answer by rfer(16322)

y=-4x-5
y+5x=3
-4x-5+5x=3
x=8
----------------
y+5*8=3
y=-37
--------------
-37+4*8=-5
-37+32=-5

Answer by ewatrrr(24785)


Hi
y+4x=-5
y+5x=3
Sets Up for a simple elimination:
-y-4x = 5 |mulitplying 1st Eq thru with -1 to eliminate y
y+5x= 3
x = 8
x = 8 and y = -37 (y +5x=3; y = -5*8 + 3)
CHECKING our Answer***
y + 4x = -5
-37 +32 = -5